How many quarts of pure (100%) antifreeze and a 50% antifreeze and 50% water mix should be combined to make 14 quarts of 70% antifreeze and 30% water?
Answer by ptaylor(2198)

Let x=amount of pure antifreeze needed
Then 14-x=amount of 50% antifreeze needed
Now we know that the amount of pure antifreeze that exists before the mixture takes place has to equal the amount of pure antifreeze that exists after the mixture takes place, sooooo:
x+0.50(14-x)=0.70*14
x+7-0.50x=9.8 simplify
0.50x=2.8
x=5.6 quarts----amount of pure antifreeze needed
CK
5.6+0.50*8.4=0.70*14
5.6+4.2=9.8
9.8=9.8
Hope this helps----ptaylor